# Personalized Topology-Informed 12-Lead ECG Electrode Localization from Incomplete Cardiac MRIs for Efficient Cardiac Digital Twins

### v1

@@ 4.7.2 Correlation between ECG Electrodes and Leads
To quantitatively assess the relationship between electrode localization error and ECG morphology misalignment (DTW), we conducted a linear regression analysis. The resulting linear coefficient $r^{2}$ was 0.492, indicating a moderately strong positive linear correlation between the localization error and morphology misalignment. To further analyze the relationship between different electrodes and leads, we calculated their Pearson correlation for the ten subjects. Figure 11 presents correlation heatmap between different electrodes and ECG leads. One can see that there exists strong positive correlation between lead I and electrode LA, as well as lead II and electrodes RA and LL. This is consistent with expectations, as lead I or II is directly related to the potential difference between two electrodes, i.e., electrodes RA and LA or electrodes RA and LL, respectively. However, it is particularly intriguing that lead V5 exhibited a notably higher dependence on the leads RA and LL. The leads V1-6 involve one exploring electrode and one Wilson’s central terminal related to electrodes RA, LA, and LL as the reference. Therefore, the correlation between leads V1-6 and electrodes could be more complex, which introduces additional variability compared to the simpler calculation between two electrodes used for leads I and II.
@@ 5 Discussion and Conclusion
In this study, we have presented a fully automatic framework for simultaneous 12-lead ECG electrode localization and 3D torso geometry reconstruction from cardiac MRI data. The proposed model was evaluated on 200 subjects from the UK Biobank, yielding better results than conventional methods ($1.24\pm 0.293$ cm vs. $1.48\pm 0.362$ cm) and significantly improved efficiency ($2$ s vs. $30$-$35$ min). This could attribute to both the topology-guided keypoint detection and the surface skeleton-assisted PCN modules, as we demonstrated in Sec. 4.5. Also, we found that random resampling of input contours had a significant impact on the performance, which might be attributed to the sparsity and incompleteness of the contours. However, this issue can be effectively addressed by increasing the number of random resampling iterations for data augmentation (see Sec. 4.4.1). The number of keypoints also affected model performance, namely they need to adequately cover the electrodes without being excessively numerous. We found that localizing electrodes LL and RL accurately was challenging, likely due to the lack of guidance from nearby landmarks (see Sec. 4.6.2). Notably, the simulated ECG morphology based on predicted electrodes exhibited good alignment with that based on ground truth (see Sec. 4.7), affirming the feasibility of employing clinically standard cardiac MRI for efficient and personalized virtual heart modeling.
@@
Nonetheless, there are three limitations of this work. Firstly, as the real electrode position is missing in the UK Biobank data, we manually localize the virtual electrodes on the reconstructed torso mesh as the ground truth. Therefore, the proposed method relies on the accuracy of the torso reconstruction, which may introduce errors if the reconstruction is not precise. Furthermore, since the rib and collarbone structure inside the torso has not been reconstructed, there may exists errors in localizing electrodes on the torso surface. Secondly, there are few studies in the literature, and no state-of-the-art 12-lead ECG electrode localization results available for comparison besides Roudijk et al. (2021). They employed 3D camera to localize 12-lead ECG electrodes for 20 atrial fibrillation patients, to reduce electrode repositioning errors during longitudinal ECG acquisition. They reported a median distance of 1.00 [0.64–1.52] cm between 3D guided repositioned electrodes, which was similar to our results ($1.24\pm 0.293$ cm). Nevertheless, we further evaluated the performance of the proposed method for in-silico ECG simulation, providing a benchmark and demonstrating its potential effectiveness in enabling an efficient CDT. Thirdly, the method has only been tested on a limited dataset for in-silico ECG simulation, and its generalizability to other datasets and populations remains to be validated. Also, this in-silico evaluation focused exclusively on the QRS complex rather than the whole ECG signal. As reported in Zappon et al. (2024), higher placement of electrodes V1 and V2 can cause noticeable T-wave inversion in ECG leads V1 and V2.
@@
In our future work, we will explore a more detailed patient-specific heart-torso anatomy, incorporating a whole heart embedded within a torso, with distinct volumes for cardiac blood pools, lungs, bones, liver, fat, and skin, to enhance the fidelity of the CDT. We will investigate the heterogeneous conductivity within the torso which presents substantial differences between compartments (including organs or distinct tissues such as bones, fat, and skin). At the same time, the electrode localization can be more accurately localized on the finer torso geometry with bone structures. We will have two experts label the virtual electrodes on the torso to assess the inter-observer variation in electrode placement. Furthermore, we will evaluate the proposed model on a larger population dataset from different centers, to investigate its generalization ability. On this larger dataset, we can perform a comprehensive sensitivity analysis to qualitatively and quantitatively investigate the impact of these uncertainties from torso and electrodes on the ECG morphology. More importantly, we will extend the in-silico ECG simulation to the entire ECG signal. We will then perform inverse inference to calibrate the simulated ECG, ensuring it aligns accurately with clinically measured ECG data. To achieve this, we will explore a computationally efficient surrogate to replace current expensive simulation model, to further speed the creation of CDT. In conclusion, this work demonstrates the feasibility of automatic electrode detection using clinically standard cardiac MRI, paving the way for future research to explore real-time CDT applications.

### v2

@@ 4.7.2 Correlation between ECG Electrodes and Leads
To quantitatively assess the relationship between electrode localization error and ECG morphology misalignment (DTW), we conducted a linear regression analysis. The resulting linear coefficient $r^{2}$ was 0.576, indicating a moderately strong positive linear correlation between the localization error and morphology misalignment. To further analyze the relationship between different electrodes and leads, we calculated their Pearson correlation for the ten subjects. Figure 11 presents the correlation heatmap between different electrodes and ECG leads. We observe a strong positive correlation between electrodes V1-V5 and their corresponding ECG leads V1-V5, with an especially high correlation for lead V3. This indicates that misplacement of electrodes in the chest area will significantly impact the morphology of their associated precordial leads, which is consistent with observations from our uncertainty study (see Supplementary Material). Additionally, electrodes LA, RA, and LL show a notable correlation with leads I and II, aligning with the findings of the uncertainty study. However, there are some unexpected correlations between certain leads and electrodes, which may indicate complex interactions due to individual anatomical differences or variations in electrode placement. Note that this analysis is based on a limited sample size of 10 subjects, which may introduce potential biases.
@@ 5 Discussion and Conclusion
In this study, we have presented a fully automatic framework for simultaneous 12-lead ECG standard torso electrode localization and 3D torso geometry reconstruction from cardiac MRI data. The proposed model was evaluated on 200 subjects from the UK Biobank, yielding better results than the conventional SSM project-based method ($1.24\pm 0.293$ cm vs. $1.48\pm 0.362$ cm) and significantly improved efficiency ($2$ s vs. $30$-$35$ min). This could attribute to both the topology-guided keypoint detection and the surface skeleton-assisted PCN modules, as we demonstrated in Sec. 4.5. Also, we found that random resampling of input contours had a significant impact on the performance, which might be attributed to the sparsity and incompleteness of the contours. However, this issue can be effectively addressed by increasing the number of random resampling iterations for data augmentation (see Sec. 4.4.1). The number of keypoints also affected model performance, namely they need to adequately cover the electrodes without being excessively numerous. We found that localizing electrodes LL and RL accurately was challenging, likely due to the lack of guidance from nearby landmarks (see Sec. 4.6.2). Notably, the simulated ECG morphology based on predicted electrodes exhibited good alignment with that based on ground truth (see Sec. 4.7), affirming the feasibility of employing clinically standard cardiac MRI for efficient and personalized virtual heart modeling.
@@
Nonetheless, there are three limitations of this work. Firstly, the proposed method aims to estimate the 12-lead ECG electrode position based on the standard protocol, assuming no significant placement errors. According to the literature, typical electrode placement errors are often within the range of 2–3 cm [17]. However, this method may not be applicable if the actual electrode placement deviates substantially from the standard protocol. Secondly, there are few related studies in the literature, and no state-of-the-art 12-lead ECG electrode localization results available for comparison besides Roudijk et al. [37]. They employed 3D camera to localize 12-lead ECG electrodes for 20 atrial fibrillation patients, to reduce electrode repositioning errors during longitudinal ECG acquisition. They reported a median distance of 1.00 [0.64–1.52] cm between 3D guided repositioned electrodes, which was similar to our results ($1.24\pm 0.293$ cm). We further evaluated the performance of the proposed method for in-silico ECG simulation, providing a benchmark and demonstrating its potential effectiveness in enabling an efficient CDT. Thirdly, the method has only been tested on a limited dataset for in-silico ECG simulation, and its generalizability to other datasets and populations remains to be validated. Also, this in-silico evaluation focused exclusively on the QRS complex rather than the whole ECG signal. As reported in Zappon et al. [49], higher placement of electrodes V1 and V2 can cause noticeable T-wave inversion in ECG leads V1 and V2.
@@
In our future work, we plan to detect the actual electrode positions by adopting the dataset collection approach described by Gillette et al. [14], which employed MRI-compatible electrodes positioned before image acquisition. Furthermore, we will explore a more detailed patient-specific heart-torso anatomy, incorporating a whole heart embedded within a torso, with distinct volumes for cardiac blood pools, lungs, bones, liver, fat, and skin, to enhance the fidelity of the CDT. We will investigate the heterogeneous conductivity within the torso which presents substantial differences between compartments (including organs or distinct tissues such as bones, fat, and skin). At the same time, the electrode localization can be more accurately localized on the finer torso geometry with bone structures. We can evaluate the proposed model on a larger population dataset from different centers, to investigate its generalization ability. On this larger dataset, we can perform a comprehensive sensitivity analysis to qualitatively and quantitatively investigate the impact of the uncertainties from torso and electrodes on the ECG morphology. These uncertainties, along with potential device artifacts in clinical ECGs, must be considered when applying predicted electrode positions to CDTs. Such factors can introduce variability in ECG morphology, potentially affecting the parameterization and accuracy of CDTs. Also, we will extend the in-silico ECG simulation to the entire ECG signal. We will then perform inverse inference to calibrate the simulated ECG, ensuring it aligns accurately with clinically measured ECG data. To achieve this, we will explore a computationally efficient surrogate to replace current simulation model, to further speed the creation of CDT. In conclusion, this work demonstrates the feasibility of automatic electrode detection using clinically standard cardiac MRI, paving the way for future research to explore real-time applications of CDTs.